OpenStudy (anonymous):

Solve for n. n/5 = 1/4

OpenStudy (anonymous):

How do you think you would start the problem \[\frac{ n }{ 5 } = \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

@domu I honestly dont know

OpenStudy (anonymous):

Okie dokie, no problem. We want to make the equation look like "n = something," the only issue we have is that we have n/5. The question now is, how do we get rid of division by 5? How do you undo division?

OpenStudy (anonymous):

how to undo division.. multiplication

OpenStudy (anonymous):

? @domu

OpenStudy (anonymous):

well continuing from domu, you multiply each side by 5 to cross the 5 out on the left side

OpenStudy (anonymous):

@domu

OpenStudy (anonymous):

so you get something like \[n =\left( \frac{ 1 }{ 4 } \right)5 \]

OpenStudy (anonymous):

@ineedhelpnow08 ok now?

OpenStudy (anonymous):

@ineedhelpnow08 is right on track, continue in this fashion

OpenStudy (anonymous):

well you convert 5 into a fraction, so 5/1 as a fraction, when you multiply, you do it straight across so u get something like \[\frac{ 1 }{ 4}\times \frac{ 5 }{ 1 } =\frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

@ineedhelpnow08 n = 1 1/4 ?

OpenStudy (anonymous):

@ineedhelpnow08 hello?

OpenStudy (anonymous):

@domu he left me can you help me

OpenStudy (anonymous):

lol sorry had to use the bathroom;......yes

OpenStudy (anonymous):

@ineedhelpnow08 so thats the answer ty!